OpenStudy (anonymous):

What is 4 over -3 multiplied by 4 over 2?

OpenStudy (anonymous):

\[\frac{4}{-3} \cdot \frac{4}{2}\]

OpenStudy (anonymous):

Multiply the numerator and denominators respectively:\[\frac{16}{-6}\]

OpenStudy (anonymous):

I worded that wrong the over should be exponents. So -3 and 2 are the exponents

OpenStudy (anonymous):

\[4^{-3} \cdot 4^{2}\] When multiplying two like bases, you add the exponents.

OpenStudy (anonymous):

So you will get \(4^{-1}\) = \(\frac{1}{4}\).

OpenStudy (anonymous):

Okay last question. Which of the transformations will always produce a congruent figure? A. a translation. And then a reflection across the x axis. B.A dialation By a scale factor of two and then a translation C. A reduction and then a 90° rotation about the origin D. A reflection and then in enlargement

OpenStudy (anonymous):

They must be the same size to be congruent, and only one of those options preserves size.

OpenStudy (anonymous):

I don't know. I'm confused

OpenStudy (anonymous):

An enlargement, dilation and reduction will all change the size.

OpenStudy (anonymous):

Translation and reflection do not affect size.

OpenStudy (anonymous):

So it would be A right?

OpenStudy (anonymous):

Yeah.